OpenStudy (anonymous):

at time t, in seconds, your velocity, v, in ft/sec, is given by v(t) = 10 - t^2 for 0≤t≤3. Use Δt = 1.5 to estimate total distance traveled during this time. find upper and lower estimates and then average the two

OpenStudy (anonymous):

i am trying to put all of the pieces together but i am not really understanding for to use Δt

OpenStudy (dumbcow):

distance = rate * time right well velocity is your rate , delta t is the time \[x(t) = v(t) * \Delta t\] your intervals for t , are 0-1.5 and 1.5 -3 for upper use t = 1.5 , t=3 for lower use t = 0, t= 1.5

OpenStudy (anonymous):

hmm okay one sec. figuring this out.

OpenStudy (anonymous):

okay now I'm just having some trouble understanding what to plug in where

OpenStudy (dumbcow):

plug in t values into velocity function , delta t is just a constant (1.5)

OpenStudy (dumbcow):

lower: \[v(0)*1.5 + v(1.5)*1.5\] upper \[v(1.5) *1.5 + v(3)*1.5\]

OpenStudy (anonymous):

okay. what exactly does Δt mean?

OpenStudy (dumbcow):

length of time .... delta just means change what this is saying is that for 1.5 sec the velocity stayed constant and the distance traveled is that velocity times time elapsed

OpenStudy (anonymous):

okay. so now it asks me to sketch a graph of velocity vs time and show the left and right rectangles. does this mean i do this on the graph of v(t)=10 - t^2??

OpenStudy (dumbcow):

yes draw the parabola , then note the points on parabola where t = 0, 1.5, 3

OpenStudy (anonymous):

thank you!!

OpenStudy (dumbcow):

yw and as you will learn , the smaller your delta t , the more accurate the estimate